Your team has allocated 30 hours per month for dealing with customer support. The time to service a customer is a normally distributed RV with a mean of 40 minutes and a variance of 49 ???2 . If during a given month there are 35 customers, what is the probability that the allocated customer support time is exceeded, (i.e. the time used by the customers is greater than the time allocated)? How many hours per month would have to be allocated to insure that for 35 customers the likelihood of exceeding the allocated time is less than or equal to 10−4 ? (40 Marks)
Answer #1

Answer:

Given,

Expected amount = 35*40 = 1400 min

Variance = 35*49 = 1715 min^2

Time = 30 hrs = 1800 min

P(X > 1800) = P(z > (1800 - 1400)/sqrt(1715))

= P(z > 9.66)

= 0

Subsequently, there is very nearly 0 likelihood of the managing time to surpass the designated esteem.

Be that as it may, on the off chance that we need the likelihood of managing time to surpass the designated an incentive to be all things considered 0.0001, at that point the z-score for the apportioned worth will be = 3.49

In this way, let X be the new dispensed time , at that point

z = (x - 1400)/sqrt(1715) = 3.49

x = 25.74 hrs